OpenStudy (anonymous):

Fine the inverse of the function. f(x)= square x-9

OpenStudy (anonymous):

hey i gave the question sweetz

OpenStudy (anonymous):

fog(x)= square x-9\[fog (x)=\sqrt{x-9}\] fog-1= for x

OpenStudy (anonymous):

y = x^2 + 9